• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gron

    Log entry time 08:17:21 on July 29, 2006

    Entry number 175898

    keyword=Owl Shift Summary

    00:00 - Started shift with no beam, due to MCC problems, they have an optices guy coming in. Due to lack of good beam during previous shift (ABU 3.3 hours) there have only been 3.8M events taken, another ~20-25M events should be taken in kinematic point #2

    1:30 - MCC says they are ready to try to send beam

    1:35 - Beam apparently good, starting runs

    2:55 - Beam Off

    6:25 - Requested beam stop in order to change kinematics

    7:00 - Started radiator thickness scan at #4 kinematics